OpenStudy (anonymous):

integral of dx/(3x^2+1)

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

it should be some type of arctan

OpenStudy (anonymous):

let me solve it one sec

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

arctan(sqrt(3)x)/sqrt(3)

OpenStudy (anonymous):

oooooo I get it...lol

OpenStudy (anonymous):

Okay! How do I give you a medal? lol

OpenStudy (anonymous):

bring three under the square, and chain rule backwards

OpenStudy (anonymous):

click good answer by my name on one of the posts

OpenStudy (anonymous):

where is the good answer? lol

OpenStudy (anonymous):

look at the grey bar with my name and picture and follow it to the right

OpenStudy (anonymous):

this should not be this difficult I'm sorry I'm new here

OpenStudy (anonymous):

haha its ok im pretty new here too. if you can't find it right now don't worry about it :)

OpenStudy (anonymous):

alright thanks! I'll still look for it

OpenStudy (anonymous):

any other questions while im here?

OpenStudy (anonymous):

hmmm yesssss

OpenStudy (anonymous):

integral x^5ln(3x)dx

OpenStudy (anonymous):

like this ? \[\int\limits_{}^{}x^{5}\ln(3x)dx\]

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

k let me work it out im geussing its by parts

OpenStudy (anonymous):

k yeah it is. Do you know integration by parts?

OpenStudy (anonymous):

ummmm kind of

OpenStudy (anonymous):

ok well I'll give you the equation: \[\int\limits_{}^{}udv=uv-\int\limits_{}^{}vdu\]

OpenStudy (anonymous):

alrighty

OpenStudy (anonymous):

Oh okay...

OpenStudy (anonymous):

so in your problem we choose ln(3x)=u, x^5=dv differentiate u and integrate dv

OpenStudy (anonymous):

okay...trying this right now...

OpenStudy (anonymous):

du=1/x v=x^6/6 so using the by parts formula now you have: \[\frac{x^{6}\ln(3x)}{6}-\int\limits_{}^{}\frac{x^{5}}{6}dx\]

OpenStudy (anonymous):

Alright...

OpenStudy (anonymous):

got that...lol

OpenStudy (anonymous):

yep so you reduce it to an integral we can do more easily like the one above

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

when I simplified I got \[\frac{x^{6}(6\ln(3x)-1)}{36}+C\]

OpenStudy (anonymous):

okay would you get...1/6x^6ln(3x) - (1/36)x^6?

OpenStudy (anonymous):

yep that is the same

OpenStudy (anonymous):

lol okay same thing I got

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

where the heck is this good anser button -.- I'm getting annoyed lol

OpenStudy (anonymous):

I'm sorry I can't find it >.<

OpenStudy (anonymous):

look at the post where I typed "no problem" it is in the upper right corner of that box

OpenStudy (anonymous):

well it should be at least haha

OpenStudy (anonymous):

it just says 0 medals for this answer o.o

OpenStudy (anonymous):

theres no button directly to the right of that in blue?

OpenStudy (anonymous):

nope o.o

OpenStudy (anonymous):

there isnt one on any of the posts? there shouldnt be one for your own, but only for others' posts

OpenStudy (anonymous):

no I don't see anything...lol

OpenStudy (anonymous):

weird, thats where it is for me.

OpenStudy (anonymous):

let me google this LOL

OpenStudy (anonymous):

i just gave you one haha

OpenStudy (anonymous):

are you a member?

OpenStudy (anonymous):

NOW I have it

OpenStudy (anonymous):

oh i geuss you need a medal to give one?

OpenStudy (anonymous):

lol I just need to refresh the page lol

OpenStudy (anonymous):

oh or maybe that too

OpenStudy (anonymous):

any more integrals?

OpenStudy (anonymous):

let me check...do you remember anything about series?

OpenStudy (anonymous):

yeah I know some about them

OpenStudy (anonymous):

We studied them a bit in complex analysis recently

OpenStudy (anonymous):

Ooo okay

OpenStudy (anonymous):

oh here is an integral...dy/dx+xy=x

OpenStudy (anonymous):

y(0)=1/2

OpenStudy (anonymous):

oh a diff equation. k give me a sec

OpenStudy (anonymous):

it is seperable. write dy/dx=x-xy=x(1-y) dy/(1-y)=xdx. Integrate both sides

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

so we get -ln(1-y)=x^2/2+C ln(1-y)=-x^2/2+C (C is some constant sign doesnt matter) 1-y=Ce^(-x^2/2) y=1+Ce^(-x^2/2) use the initital condition to find C y(0)=1/2 1/2=1+C C=-1/2 y=1-(1/2)e^(-x^2/2)

OpenStudy (anonymous):

Oooo so you want to put one side over the other? and then integrate

OpenStudy (anonymous):

you want to get the y terms with dy and x terms with dx so you can integrate both sides

OpenStudy (anonymous):

Ooo okay

OpenStudy (anonymous):

if you can do this it is called a seperable equation

OpenStudy (anonymous):

that should be enough! :) thank you very much

OpenStudy (anonymous):

lol this is a review...but I don't remember lol

OpenStudy (anonymous):

there is a lot of material in calculus

OpenStudy (anonymous):

do you want to do one series or good for now?

OpenStudy (anonymous):

Ummmm good for now :)

OpenStudy (anonymous):

k talk to you later

OpenStudy (anonymous):

Thank you very much! Tuly appreciate it

OpenStudy (anonymous):

you're welcome